solve a uniformly capillary tube closed at one end contain dry air trapped by a thread of mecury 8.5*10m long. when the tube was held horizontally the length of the air column was 5.0^-2m when it was held vertically with the closed end downward length was 4.5*10^-2m determine the value of atmosphere pressure ( g=10m/s^-2) density of mecury 1.36*104kgm^-3


1
Expert's answer
2021-10-04T19:36:05-0400

"P_1V_1=P_2V_2" ..........(i)

"V_2=\frac{P_1V_1}{P_2}...........(ii)"

"P_1V_1=h_1(h_m+atm)"

"P_2=atm"

"V_2(horizntal)=5\times 10^{-2}m"

"h_1(closed)=4.5\times 10^{-2}m"

"h_m=8.5\times 10^{-2}m"

substituting these values in equation (ii) we have

"V_2=\frac{P_1V_1}{P_2}=\frac{h_1(h_m+atm)}{atm}"

"5\times 10^{-2}=\frac{4.5\times 10^{-2}(8.5\times 10^{-2} +atm)}{atm}"

"0.05atm=0.003825+0.045atm"

"0.005atm=0.003825"

"atm=0.765mHg=76.5cmHg"